1.a)Draw an A.D/A.S graph showing inflationary gap.

b)Draw an A.D/A.S graph showing a recessionary gap.

c)Using the same graphs you have drawn in A and B, show what the long run equilibrium position would be in each case if the government did nothing(i.e.,let the economy self-adjust)

2.Explain how savings function is derived from the consumption function, and how the saving schedule and graph of the saving function are constructed


1
Expert's answer
2021-05-06T13:34:44-0400

1.

a) Inflationary Gap is shown in the diagram using AD-AS model. 





Here, the actual AD is more than the potential AD, Aggregate demand is greater than Aggregate Supply. Yf-Y is the inflationary gap. 


b) Recessionary Gap is also called deflationary gap which is shown in the diagram below using AD-AS model.



Here, the actual Aggregate demand (AD') is less than potential Aggregate demand(AD) . Thus, AS> AD' leading to recessionary gap of Y-Yf. 


c) When there is no government intervention, the economy adopts the self correcting measures.

CASE I : Inflationary Gap

In the long run, the market will eliminate the inflationary gap using the self adjustment policy as there is shortage in the economy. It is done by increase in wages which will lead to shift in the short run aggregate supply curve. There will be higher wages offered which in turn will decrease the Aggregate supply in the short run leading to leftward shift in the curve from AS to AS'.



The new equilibrium is achieved at AS' and AD2, which is the short run curve which also coincides with long run AS curve at E*. Inflationary gap is eliminated with higher wages and decrease in aggregate supply.


CASE II : Recessionary Gap

In the long run, the market with recessionary gap will eliminate the recessionary gap with the help of self adjustment policy. As the output level is below the full employment level, lower wages will help to eliminate the gap by increasing the short run aggregate supply curve from AS to AS'. This will bring the economy to full employment level of output. 




The new equilibrium is achieved at E* where the actual AD' matches short run aggregate supply curve which is also the point of intersection of LRAS. Thus full employment level of output is achieved through self adjustment by lower wages and increased aggregate supply in long run. 


2.

Saving Function is simply the will of the households to hoard a neighborhood of their total income .

Symbolically, the functional relation between saving and income are often defined as"S= f(Y)."

we know,

"Y= C + S;"

"Thus, S= Y-C;"


Where;

Y= Income

S= Saving

C= Consumption

 




The equation explains that the remaining amount after the deduction of total expenditure from total income is saving. Thus, saving is that a part of income which isn't spent on consumption